OpenStudy (anonymous):

let f(x)=x^3-x^2+x. show that there is a numbr c such that f(c)=10

OpenStudy (anonymous):

I think you would do 10=x^3-x^2+x

OpenStudy (anonymous):

Thanks. But wat would you do then

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

i don't believe it is true, since you have a cubic function which goes from \(-\infty\) to \(\infty\)

OpenStudy (anonymous):

oooh i see doh it says show that there IS a number no problem

OpenStudy (anonymous):

any cubic function has range all real numbers, so of course it is true

OpenStudy (anonymous):

how do you know its true? could you like, solve the problem for me so I could see and learn frm how you did it?

OpenStudy (anonymous):

in your case as \(x\to -\infty\) you know \(f(x)=x^3-x^2+x\to -\infty\) and as \(x\to \infty\) you have \( f(x)=x^3-x^2+x\to \infty\) and since all polynomials are continuous, it must be equal o ten somewhere

OpenStudy (anonymous):

another way you could to it is pick some number for \(a\) so that \(f(a)<10\) and some number \(b\) so that \(f(b)>10\) and then say "since \(f\) is a polynomial it is continuous, so it must take on all values between \(f(a)\) and \(f(b)\)

OpenStudy (anonymous):

the point is a cubic function is onto so it takes on every value